OpenStudy (andromeda992):

Please help! A quantity of gas has a volume of 0.20 cubic meter and an absolute temperature of 333 degrees kelvin. When the temperature of the gas is raised to 533 degrees kelvin, what is the new volume of the gas? (Assume that there's no change in pressure.) A. 0.0006 m^3 B. 0.2333 m^3 C. 0.2146 m^3 D. 0.3198 m^3

OpenStudy (quennlee):

Use the Charles' Law. Are you familiar with that?

OpenStudy (quennlee):

V1/T1 = V2/T2

OpenStudy (andromeda992):

I tried that but I got a completely different answer then anything listed

OpenStudy (andromeda992):

\[\frac{ 0.20}{ 333K }=\frac{ V2 }{ 533 }\]

OpenStudy (quennlee):

But the problem said that it has an absolute temperature and not only that.. the problem also said that there's no change in pressure..

OpenStudy (quennlee):

You can refer on this site: http://scienceprimer.com/charles-law It might help you to fully understand the question. :)

OpenStudy (quennlee):

TAKE NOTE: Not all choices are the required answer. :)

OpenStudy (quennlee):

Thanks for the medal. Hehe. ^_^

OpenStudy (andromeda992):

you're welcome